OpenStudy (anonymous):

2ln3+ln(x+1)=0 solve for x

OpenStudy (amoodarya):

2ln3+ln(x+1)=0 ln9+ln(x+1)=0 ln(9(x+1))=0 so 9x+9=e^0=1 so 9x=1-9 9x=-8 x=-8/9

OpenStudy (tkhunny):

The very first thing you OUGHT to do is consider the Domain. A quick look at the ORIGINAL problem statement suggests \(x+1 \gt 0\;or\;x \gt -1\). This sort of thing prevents us from believing entirely inappropriate results. In this case, \(-8/9 \gt -1\), so this seems appropriate. Had we concluded x = -11/9, we would have discarded it summarily.